The incubation time for chicks is normally distributed with mean 21 days and standard deviation approximately 1 day. If 1000 eggs are being incubated, how many chicks do we expect will hatch?
a) in 19 to 23 days?
b) in 22 days or more?
c) in 21 days or fewer?
d) in 18 to 21 days?
e) in 19 days or fewer
f) in 19 to 24 days?
g) in 18 to 24 days?
h) in 18 days or more?


Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

mean = 21 days
standfard deviation = 1 day

total of 1000 eggs are being incubated.

how many checks are expected to hatch in:

a) in 19 to 23 days?
.954 * 1000 = 954

b) in 22 days or more?
.159 * 1000 = 159

c) in 21 days or fewer?
.5 * 1000 = 500

d) in 18 to 21 days?
.499 * 1000 = 499

e) in 19 days or fewer
.023 * 1000 = 23

f) in 19 to 24 days?
.976 * 1000 = 976

g) in 18 to 24 days?
.997 * 1000 = 997

h) in 18 days or more?
.999 * 1000 = 999



Answer by ikleyn(52878)

.


        Tutor @Theo in his post provided the answers, but did not provide the method.


Use online free of charge calculator for normal distribution at this web-site

https://onlinestatbook.com/2/calculators/normal_dist.html

This calculator has convenient, user-friendly interface, intuitively understandable
from the first glance. It ideally fits for quick calculations,
and it ideally fits to be an excellent teacher/instructor for such calculations.

Find the answers in the post by @Theo to this problem.